Will runs a dog grooming service, and has asked you to analyze the current 'line situation'. He has gathered data and calculated that he can groom a dog in about 10 minutes, based on a negative exponential distribution. He also noted that about 3 dogs arrive to his shop per hour, during all operating hours, based on a Poisson distribution.

How long, on average, will a dog wait for grooming? ["10 minutes", "1 minute", "16.67 minutes", "20 minutes"]      

In total, how long will a dog spend in the shop waiting and being groomed? ["26.67 minutes", "20 minutes", "16.67 minutes", "10 minutes"]      

Will's shop is fairly small, and he only has space for 3 dogs to be in the shop before it gets crowded. What is the probability that there will be more than 3 dogs at any time? ["3.125%", "6.25%", "1.5625%", "12.5%"]      

Will wants to reduce the wait time, and is considering two options. The first would be to hire a helper, who would assist Will, and reduce the grooming time from 10 minutes to 8 minutes. (NOTE: The helper only helps, and does not work alone.)

What would the wait time be with a helper? ["3.33 minutes", "1.33 minutes", "13.33 minutes", "5.33 minutes"]      

The second option is to rent an Auto-Groomer, a new machine that would allow Will to groom a dog in a constant 9 minutes.

What would the wait time be with Auto-Groomer? ["3.68 minutes", "4.67 minutes", "3.33 minutes", "5.33 minutes"]      

Assuming both options cost the same amount and all else is equal, which should Will choose to deliver the shortest wait time? ["Choose the helper", "Choose either; wait times are equal", "Choose the Auto-Groomer"]      
Answer #1

Average rate of arrivals per hour , = 1/4= 3 Dog(s) per hour 0.3333 hrs = 20 mins Time between arrival of Dog(s)(A) = Average

For auto groomer option1/A=3 Dog(s) per hour Average rate of arrivals per hour ,2 Time between arrival of Dog(s)(A) Average rate of service per hour

As it is clear, the costs being the same for both the options we need to go with the Helper option because of the waiting time in the queue in case of a helper than the other option. This will also provide the opportunity for serving more dogs than the option 2